### v1

# Parallel Batch Scheduling With Incompatible Job Families Via Constraint Programming

Abstract
This paper addresses the incompatible case of parallel batch scheduling, where compatible jobs belong to the same family, and jobs from different families cannot be processed together in the same batch. Existing constraint programming (CP) models for this problem fail to synchronize the processing of the jobs within their batch, resulting in batch interruptions. In the context of the diffusion area in the semiconductor manufacturing process, these interrupted solutions would disrupt the thermal stability required for a uniform dopant distribution on the wafers. This paper proposes three new CP models that directly tackle these interruptions in the formulation, including two adaptions of existing models and a novel Redundant Synchronized (RS) model. These existing and novel models are compared on standard test cases, demonstrating the superiority of the RS model in finding optimal or near-optimal solutions quickly.
Index Terms:
Scheduling, Parallel Batching, Incompatible Job Families, Constraint Programming, Synchronization.
I Introduction
In the current and highly competitive landscape of the manufacturing industry, companies are under growing pressure to minimize production costs and reduce cycle times. A crucial approach to achieving these goals is performing certain manufacturing processes in batches [1, 2]. In parallel batching (p-batch), jobs inside a batch are processed in parallel at the same time, allowing to reduce their cycle times [2, 3, 4].
P-batch is a common problem that appears in many contexts such as metal processing [5], truck delivery [6], automobile gear manufacturing [7], additive manufacturing (AM) [8, 9, 10], composite manufacturing [11], milk processing [12], and many more [2, 13]. Nonetheless, it has been mostly studied in the semiconductor manufacturing (SM) process, which is the most complex manufacturing process in the world and consists of two stages [1, 2, 3]. The first stage is the front-end, where the integrated circuits (ICs) are built on the silicon wafers in a wafer fabrication facility (wafer fab). The second stage is the back-end, where the individual ICs are assembled (A) and tested (T) before final packaging in an AT fab. [2, 3].
In the back-end, p-batching has been typically studied in the burn-in operations of the testing phase, where the individual ICs are subject to thermal stress over an extended period to detect latent defects [2, 14]. The test of an individual IC is a job that requires a minimum processing time, predefined ahead of time. Due to the prolonged processing times, jobs are batched to be put in the oven, and the processing time of the batch is the longest of the jobs in it. In this setting, jobs can remain in the oven beyond their required time but cannot be removed before the batch is completed [15, 16]. This scheduling problem belongs to the compatible case of p-batch, where all the jobs are compatible and can therefore be batched together [2].
A more complex p-batch problem arises in the front-end of the SM process, in the diffusion area of the wafer fab, where dopants are introduced into the silicon wafers at high temperatures to alter their electrical properties. This is achieved using diffusion furnaces that operate in batches, applying constant heat on the wafers for long periods, typically 10-12 hours [17]. These prolonged processing times are essential to ensure the thermal stability required for an even distribution of dopants on the wafers [18]. Due to the chemical nature of the process, not all the jobs can be batched together. Instead, incompatible job families are defined by the distinct “recipes” of the jobs, which dictate unique dopants and thermal conditions [2]. Jobs with the same recipe are said to be compatible and therefore grouped into the same family; jobs in the same family also have the same processing time [1, 19]. Jobs of different families are said to be incompatible and cannot be processed together in the same batch, as this may result in unwanted cross-contamination in the chemical process. Hence, the scheduling problem in the diffusion area belongs to the incompatible case of p-batch, where the batches can only contain jobs of the same family [2, 18]. Furthermore, due to the risk of burning the wafers or damaging the chemical process, jobs cannot remain in the furnace beyond their processing time. Moreover, the processing of a batch cannot be interrupted to include or remove jobs mid-process, as this would disrupt the thermal stability and jeopardize the uniformity of the diffusion [18]. Thus, jobs inside a batch must be synchronized, i.e., their processing must start and end together. This synchronization of jobs within their batch is critical for the diffusion but is typically neglected in traditional p-batch models, like the ones for the compatible case of p-batch in the burn-in operations.
Existing exact methods in the literature for p-batching include mixed-integer programming (MIP) [6, 18] and Constraint Programming (CP) models [8, 18]. MIP models do not scale well and are outperformed by the CP approaches [2, 18]. However, the existing CP models fail to synchronize the jobs within their batch, allowing interruptions mid-process. This paper remedies this shortcoming by proposing three new CP models that directly tackle these interruptions in their formulation, and explores the impact of breaking symmetries in the solution. Furthermore, the paper conducts thorough computational experiments that demonstrate the benefits of the proposed models over those in the literature.
The remainder of this paper is organized as follows. Section II presents a literature review on p-batching. Section III clearly outlines our contributions. Section IV formally presents the description of the problem we address. Section V presents an existing CP model that allows interruptions. Section VI presents our proposed CP models that remedy them. Section VII presents our computational experiments and their results. Finally, Section VIII presents the conclusions and outlines future lines of research.
II Literature Review
The scheduling literature distinguishes two types of p-batch: compatible and incompatible p-batch [2, 1, 3]. In the compatible case, all the jobs are compatible and can therefore be batched together. In contrast, in the incompatible case, jobs are grouped into families, compatible jobs belong to the same family, and jobs of different families cannot be processed together in the same batch [3, 2, 4].
The standard classification scheme for scheduling problems provided by Graham et al. [20] is $\alpha|\beta|\gamma$, where $\alpha$ denotes the scheduling environment, $\beta$ describes the job and family characteristics and any restrictions, and $\gamma$ defines the objective function. This section focuses on relevant articles that deal with multiple parallel machines (i.e., $\alpha=P$). For a more comprehensive overview of single-machine p-batch, interested readers are referred to the recent survey by Fowler and Mönch [2]. For $\beta$, the focus is on articles that deal with multiple incompatible families (i.e., $P|p-batch,incompatible|\gamma$), although key articles with a single compatible family (i.e., $P|p-batch|\gamma$) are also considered, as the paper extends some of their concepts to the incompatible case. The review also includes some articles that consider non-identical job sizes ($s_{j}$) and release times ($r_{j}$). For $\gamma$, the reviewed objective functions primarily fall into three categories: (i) utilization objectives such as minimizing the makespan; (ii) cycle time objectives such as minimizing the total completion time (TCT) and its weighted counterpart (TWCT); and (iii) on-time delivery objectives such as minimizing earliness, lateness, tardiness, number of tardy jobs, including their weighted variants [2]. However, since the diffusion area is usually a bottleneck in the wafer fab, fab leaders often focus on the cycle time of the jobs to increase the effective movement of products [19]. Thus, the focus is on minimizing the TWCT, which specifically targets the cycle time performance [2].
Most of the articles that consider multiple parallel machines usually focus on a single family (i.e., $P|p-batch|\gamma$) [2]. This problem is NP-hard [21, 2] and multiple heuristic algorithms [22, 23, 24] as well as meta-heuristics [25, 26, 27] have been proposed. Some exact approaches include mixed-integer linear programming (MILP) models [16] and efficient arc-flow formulations [28, 29, 30, 5]. Recently, CP has been used for this problem in contexts different from SM, such as the composite manufacturing industry [11], and AM processes where parts with irregular shapes are 3D printed [8, 9, 10]. Nascimento et al. [8, 9] presented an Assign-and-Schedule (AS) CP model that uses binary variables for the assignment of the parts to printing batches and CP scheduling logic to sequence the batches on the 3D printers. Their assignment logic already locates the parts inside the batch, so they do not explicitly include capacity constraints. However, they only consider compatible jobs (i.e., one family). This paper extends their formulation to handle multiple incompatible families, and include additional constraints to handle the batch size requirements.
In another recent development in the AM industry, Cakici et al. [10] presented a synchronized CP model that replaces binary variables with interval variables, resulting in improved performance over the AS formulation. However, they consider only a single family of compatible jobs. Furthermore, their model leverages optional interval variables in combination with a global synchronization constraint to ensure that jobs within the same batch start and finish together. While this approach is computationally efficient, it can lead to synchronization inconsistencies due to the optional nature of the variables involved in the synchronization process. Specifically, their model does not explicitly enforce the presence of the required batches when synchronizing jobs, which may result in unintended outcomes. This paper addresses this limitation by introducing a constraint that guarantees the presence of a batch whenever a job is assigned to it, thereby ensuring consistent synchronization across all jobs and batches. Moreover, this paper extends their model to handle multiple incompatible job families and impose additional constraints to accommodate batch size requirements.
The literature that deals with multiple parallel machines and incompatible job families (i.e., $P|p-batch,incompatible|\gamma$) is rather scarce [2]. Since this problem is NP-hard [4], the literature is dominated by heuristics and meta-heuristics [4, 31, 32, 33, 34, 35, 19]. However, exact methods have also been proposed, such as DP [19] and MIPs [6, 18]. Cakici et al. [6] proposed a time-indexed (TI) MIP model inspired by the truck delivery problem. They also presented two variable-neighborhood search (VNS) heuristics that outperform the MIP model. Later, Ham et al. [18] presented another MIP formulation based on positional assignment (PA) variables in the context of the diffusion area. They also proposed a CP model that outperforms both the PA and the TI MIP models, as well as the VNS heuristics. However, this CP model fails to synchronize the jobs within their batch, allowing job interruptions mid-process. This paper addresses this issue and proposes a set of synchronized CP models that tackle the interruptions directly in the formulation.
III Contributions
As evidenced in the previous section, most of the existing CP models for p-batch only deal with a single compatible family [8, 9, 10]. Among these, the only model that synchronizes the jobs does not enforce the presence of a batch when jobs are assigned to it [10], resulting in an inconsistent schedule. Moreover, the existing CP model that considers multiple incompatible families fails to synchronize the jobs within their batch, allowing interruptions mid-process [18].
This paper addresses these gaps in the literature by proposing three new CP models that consider multiple incompatible job families and directly tackle the interruptions in their formulation while synchronizing the jobs within their batch: (i) An Assign-and-Schedule (AS) model inspired by Nascimento et al. [8, 9] that handles the assignment of jobs to batches with binary variables, and the sequencing of batches on machines using CP scheduling logic; (ii) A Synchronized (S) model inspired by Cakici et al. [10] that replaces the binary assignment with CP scheduling logic, and guarantees the correct job synchronization by ensuring the “presence of batches” that have assigned jobs; (iii) and a new Redundant Synchronized (RS) model that includes redundant variables and constraints that help the search process and improves computational and quality performance. Furthermore, the paper explores the impact of symmetry-breaking constraints in the search process. The computational experiments demonstrate the superiority of the RS model over existing models.
IV Problem Description
Let $\mathcal{J}$ be the set of jobs, $\mathcal{F}$ be the set of incompatible job families, and $\mathcal{M}$ be the set of machines. All jobs can be scheduled on all machines. Compatible jobs belong to the same family. Hence, let $f_{j}\in\mathcal{F}$ be the family of job $j\in\mathcal{J}$, and $\mathcal{J}_{f}=\set{j\in\mathcal{J}:f_{j}=f}$ be the subset of jobs that belong to family $f\in\mathcal{F}$. All the jobs of the same family have the same processing time, thence, let $p^{f}$ be the processing time of family $f$ and let $p_{j}=p^{f}$ be the processing time of job $j\in\mathcal{J}_{f}$. Additionally, each job $j$ has a size $s_{j}$ (i.e., number of parts), a weight $\omega_{j}$ (i.e., priority), and a release time $r_{j}$. Each machine $m\in\mathcal{M}$ can simultaneously process multiple jobs of the same family $f\in\mathcal{F}$ in batches whose size must be between $l_{f}$ and $u_{f}$, as evidenced in the instances of the SM testbed 2020 [17]. Jobs in the same batch have to start and end together in a synchronized fashion, and jobs of different families cannot be processed together in the same batch. The goal is to minimize the total weighted completion time (TWCT) of the jobs. In the classification by Graham et al. [20], this problem is $P|s_{j},r_{j},p-batch,incompatible,synchronized|TWCT$.
V The Existing CP Model
CP is a powerful method for solving combinatorial problems by defining constraints that a solution must satisfy. In CP, variables are assigned values from their domains, and the system works to ensure that all constraints are met. The search process explores the solution space in the form of a tree, where fixing the values of variables at each node branches out potential solutions. A key aspect of CP is the use of constraint propagation, which reduces the search space by pruning values from variable domains that cannot satisfy the constraints. Backtracking occurs when a branch leads to an invalid solution, allowing the search to revert to an earlier state and explore different variable assignments [36].
Scheduling is one of the most successful application areas of CP. It leverages interval variables to represent tasks or operations over time. These interval variables encapsulate three key components: the start time, the end time, and the presence status of the interval (indicating whether the task is executed). The start time and the end time define the size of the interval. The presence attribute allows CP to model both mandatory and optional activities within a schedule. In particular, it plays a crucial role in handling optional tasks and alternative resource allocations. Interval sequence variables allow to model sequences of tasks. CP also uses cumulative functions to model the evolution of resource usage over time, ensuring that resource capacities are respected by aggregating the resource demands of all overlapping tasks. These cumulative functions are crucial in preventing over-utilization of resources like machines, personnel, or energy. The efficiency of the search process is further enhanced by global constraints that exploit the structure of the problem to propagate information and prune variable domains efficiently. Together, CP and its tools for constraint propagation, domain pruning, backtracking, and cumulative functions provide a structured and efficient approach for solving complex scheduling problems [36, 37]. The interested reader on an in-depth CP overview is referred to the well-known book by Van Beek et al. [36], and the paper by Laborie et al. [37] about IBM ILOG CP Optimizer.
Model 1 presents the CP model proposed by Ham et al. [18] to schedule incompatible jobs in the diffusion area. Equation (1) defines an interval variable $x_{j}$ of the form $[a,a+p_{j})$ (i.e., it has a duration of exactly $p_{j}$ units of time), which represents job $j\in\mathcal{J}$. Equation (1) defines an optional interval variable $x_{jm}$, also of size $p_{j}$, that represents the option of processing job $j$ on machine $m\in\mathcal{M}$. The presence of the interval $x_{jm}$ is optional since it is allowed to take the value $\perp$, which indicates its absence in the solution. These optional interval variables allow the model to decide which machine should process the jobs. Equation (1) defines a function ${\mathfrak{f}}_{m}$, which is the unique state of the machine $m$ at every point in time. Equation (1) defines a cumulative function ${\mathscr{l}}_{mf}$, which represents the load on machine $m$ of jobs of the family $f\in\mathcal{F}$. It uses the pulse function, which receives two inputs: An interval $x_{jm}$ and an integer value $s_{j}$. This pulse function tells the cumulative function to increment its value in $s_{j}$ units during the interval $x_{jm}$, if present. Hence, the function ${\mathscr{l}}_{mf}$ accumulates the sizes of overlapping jobs of family $f\in\mathcal{F}$ on machine $m$ at every point in time.
Model 1 Non-synchronized model by Ham et al. [18]
Variables:
$x_{j}$$\in\{[a,a+p_{j}):a\in\mathbb{Z}\},$ $\forall\leavevmode\nobreak\ j\in\mathcal{J};$ $(1)$
$x_{jm}$$\in\{[a,a+p_{j}):a\in\mathbb{Z}\}\cup\{\perp\},$ $\forall\leavevmode\nobreak\ j\in\mathcal{J},m\in\mathcal{M};$ $(2)$
Functions:
[…]
${\mathscr{l}}_{mf}$ $=\sum_{j\in\mathcal{J}_{f}}\text{pulse}\left(x_{jm},s_{j}\right),$ $\forall\leavevmode\nobreak\ m\in\mathcal{M},f\in\mathcal{F}.$ $(4)$
Formulation:
minimize $\text{minimize\leavevmode\nobreak\ }\sum_{j\in\mathcal{J}}\omega_{j}\cdot\texttt{endOf}\left(x_{j}\right),$ $(5)$
subject to,
$\forall\leavevmode\nobreak\ j\in\mathcal{J}:$ (6)
           alternative $\left(x_{j},\set{x_{jm}}_{m\in\mathcal{M}}\right);$
$\forall\leavevmode\nobreak\ m\in\mathcal{M},j\in\mathcal{J}:$ (7)
alwaysEqual $\left({\mathfrak{f}}_{m},x_{jm},f_{j}\right);$
$\forall\leavevmode\nobreak\ m\in\mathcal{M},f\in\mathcal{F},j\in\mathcal{J}_{f}:$ (8)
alwaysIn $\left({\mathscr{l}}_{mf},x_{jm},l_{f},u_{f}\right).$
Objective function (1) minimizes the TWCT. Constraints (1) use the alternative global constraint, which receives two inputs: an interval variable $v=x_{j}$, and a set of optional interval variables $B=\set{x_{jm}}_{m\in\mathcal{M}}$. This global constraint ensures that, if the interval variable $v$ is present, then only one of the optional interval variables in $B$ gets selected to be present, and $v$ gets synchronized with the selected interval in $B$ [38]. Hence, constraints (1) enforce each job to be processed on exactly one machine. Constraints (1) use the global constraint alwaysEqual, which receives three inputs: A state function $h={\mathfrak{f}}_{m}$, an interval variable $v=x_{jm}$, and an integer value $a$. This global constraint ensures that the state function $h=a$ during the interval $v$, if present [38]. Hence, constraints (1) ensure that jobs of only one family are processed at any time on the machines. Finally, constraints (1) use the global constraint alwaysIn, which receives four inputs: A cumulative function $g={\mathscr{l}}_{mf}$, an interval variable $v=x_{jm}$, and integer values $a=l_{f}$ and $b=u_{f}$. This global constraint ensures that $a\leq g\leq b$ during the interval $v$, if present [38]. Hence, constraints (1) guarantee that the load on the machines is within the required batch size limits of the family being processed.
V-A Shortcomings of the Existing CP Model
Model 1 clearly selects one machine for each job, guarantees that only jobs of one family are processed at the same time on each machine, and ensures that the load on each machine is within the batch size limits of the family being processed on it. Nonetheless, the model never ensures that the jobs being processed on the machines at any point in time start and end together in a synchronized fashion. Model 1 is called Non-synchronized (NS), as it allows job interruptions mid-process.
Consider the scenario in Table I as an illustrative example with four jobs, one family, and one machine. Assume that the size of each batch must be between $l_{1}=50$ and $u_{1}=100$. Figure 1 shows the Gantt chart of the optimal solution achieved by the NS model. The thick blue dotted line shows the load on the machine with jobs of family $1$. This solution has a TWCT $=1,550$. The state of the machine is the family of the jobs being processed, representing a batch of jobs of family $1$. However, the jobs are not synchronized, allowing two batch interruptions and four job interruptions. At time 12, jobs 1 and 3 are interrupted to include jobs 2 and 4. At time 15, jobs 2 and 4 are interrupted to remove jobs 1 and 3. If these interruptions happen in the diffusion area, the thermal stability required by the wafers would be damaged.
Table I: Information of the illustrative example
Job $j$ Size $s_{j}$ Weight $\omega_{j}$ Release time $r_{j}$ Family $f_{j}$ Processing time $p_{j}$
1 25 10 3 1 10
2 25 10 11 1 10
3 25 20 5 1 10
4 25 40 12 1 10
Figure 1: Optimal solution of the Non-synchronized model
V-B Feasibility Restoration
The NS model allows job interruptions, making the solutions infeasible to the problem. To remedy this, after retrieving the solution of the NS model, this paper proposes to conduct a Feasibility Restoration (FR) step as follows: the scheduled jobs on each machine are sorted by their start time, and the FR iterates through them while keeping track which jobs start, which jobs are currently in process, and which jobs are completed. If there are jobs currently in process when a new job starts, FR either pulls the interrupted jobs to start with the new ones, or pushes the starting jobs right after the jobs that are currently in process (and shifts all the future jobs accordingly to maintain the sequence). This local decision of pulling or pushing jobs is made greedily while ensuring feasibility of the batch size limits and compatibility of batched jobs. By the end of the process, the solution is guaranteed to be feasible without interruptions. Nonetheless, no guarantee on the quality of the resulting solution can be made. To remedy this, Section VI proposes three new CP models that directly tackle interruptions in their formulations.
VI The Proposed CP Models
This section presents the three proposed CP models that directly handle the interruptions in their formulation. These models make explicit use of the ordered set of possible batches =$\mathcal{B}=\set{1,2,3\ldots}$. The family $f^{b}\in\mathcal{F}$ of each possible batch $b\in\mathcal{B}$ is defined a priori. In addition, the model only allows jobs $\mathcal{J}_{f^{b}}$ of family $f^{b}$ to be processed on batch $b$. Hence, this set can be further partitioned into the mutually exclusive sets $\mathcal{B}_{f}=\set{b\in\mathcal{B}:f^{b}=f}\subset\mathcal{B}$ of possible batches for each family $f\in\mathcal{F}$. In this way, the first $|\mathcal{B}_{1}|$ elements of $\mathcal{B}$ correspond to the possible batches where jobs of family $1$ can be processed; the next $|\mathcal{B}_{2}|$ elements to the possible batches for jobs of the family $2$, and so on. The number of possible batches of each family $|\mathcal{B}_{f}|$ is upper bounded by the number of jobs that belong to the family $|\mathcal{J}_{f}|$. On the other hand, assuming $l_{f}>0$, this number is also upper bounded by $\lceil\frac{N_{f}}{l_{f}}\rceil$, where $N_{f}=\sum_{j\in\mathcal{J}_{f}}s_{j}$ is the aggregated size of all the jobs of family $f$. Consequently, we set $|\mathcal{B}_{f}|=\min\set{}{\mathcal{J}_{f}|,\lceil\frac{N_{f}}{l_{f}}\rceil}$.
VI-A The Assign-and-Schedule Model
Model 2 The Assign-and-Schedule Model Inspired by [8]
Variables:
$z_{jb}$$\in\set{0,1},$ $\forall\leavevmode\nobreak\ j\in\mathcal{J},b\in\mathcal{B}_{f_{j}};$ $(9)$
$y_{b}$$\in\{[a,a+p^{f^{b}}):a\in\mathbb{Z}\}\cup\{\perp\},$ $\forall\leavevmode\nobreak\ b\in\mathcal{B};$ $(10)$
$y_{bm}$$\in\{[a,a+p^{f^{b}}):a\in\mathbb{Z}\}\cup\{\perp\},$ $\forall\leavevmode\nobreak\ b\in\mathcal{B},m\in\mathcal{M};$ $(11)$
$\varphi_{m}$ $\in\text{Perm}\left(\set{y_{bm}}_{b\in\mathcal{B}}\right),$ $\forall\leavevmode\nobreak\ m\in\mathcal{M}.$ $(12)$
Formulation:
minimize $\text{minimize\leavevmode\nobreak\ }\sum_{j\in\mathcal{J}}\omega_{j}\cdot\sum_{b\in\mathcal{B}_{f_{j}}}\texttt{endOf}\left(y_{b}\right)\cdot z_{jb},$ $(13)$
subject to,
$\forall\leavevmode\nobreak\ j\in\mathcal{J}:$ (14)
$\sum_{b\in\mathcal{B}_{f_{j}}}z_{jb}\nobreak\leavevmode$ $=1;$
$\forall\leavevmode\nobreak\ b\in\mathcal{B}:$
$\sum_{j\in\mathcal{J}_{f^{b}}}z_{jb}\nobreak\leavevmode$ $\geq\texttt{presenceOf}\left(y_{b}\right);$ $(15)$
$\sum_{j\in\mathcal{J}_{f^{b}}}z_{jb}\nobreak\leavevmode$ $\leq|\mathcal{J}_{f^{b}}|\cdot\texttt{presenceOf}\left(y_{b}\right);$ $(16)$
$\forall\leavevmode\nobreak\ b\in\mathcal{B}:$
$\sum_{j\in\mathcal{J}_{f^{b}}}s_{j}\cdot z_{jb}\nobreak\leavevmode$ $\geq l_{f^{b}}\cdot\texttt{presenceOf}\left(y_{b}\right);$ $(17)$
$\sum_{j\in\mathcal{J}_{f^{b}}}s_{j}\cdot z_{jb}\nobreak\leavevmode$ $\leq u_{f^{b}}\cdot\texttt{presenceOf}\left(y_{b}\right);$ $(18)$
$\forall\leavevmode\nobreak\ b\in\mathcal{B},j\in\mathcal{J}_{f^{b}}:$ (19)
$\texttt{startOf}\left(y_{b}\right)$ $\geq r_{j}\cdot z_{jb};$
$\forall\leavevmode\nobreak\ b\in\mathcal{B}:$ (20)
alternative $\left(y_{b},\set{y_{bm}}_{m\in\mathcal{M}}\right);$
$\forall\leavevmode\nobreak\ m\in\mathcal{M}:$ (21)
noOverlap $\left(\varphi_{m}\right).$
Model 2 presents the Assign-and-Schedule (AS) model. It is based on the CP model by Nascimento et al. [8] for AM processes, which deals with a single compatible family and does not include capacity constraints. This paper extends their formulation to handle multiple incompatible families and include the necessary constraints to ensure the batch sizes. The model name originates from its clear distinction between assignment and scheduling decisions. While the job assignment to batches is handled with binary variables, the batch sequencing is handled with CP scheduling logic.
Equation (2) defines a binary variable $z_{jb}$ that takes the value of 1 if job $j\in\mathcal{J}$ is assigned to batch $b\in\mathcal{B}_{f_{j}}$, or 0 otherwise. Equation (2) defines an optional interval variable $y_{b}$ of size $p^{f^{b}}$ that represents batch $b\in\mathcal{B}$. It is optional because a batch with no jobs assigned should not be present in the solution. Equation (2) defines an optional interval variable $y_{bm}$, also of size $p^{f^{b}}$, that represents batch $b$ being processed on machine $m\in\mathcal{M}$. Equation (2) defines an interval sequence variable $\varphi_{m}$, which represents the sequence of batches scheduled on machine $m$. This sequence variable can be any permutation of the batch intervals $\set{y_{bm}}_{b\in\mathcal{B}}$ on the machine $m$ present in the solution.
Objective function (2) minimizes the TWCT. Constraints (2) assign each job to exactly one possible batch. Constraints (2) deactivate the presence of the batches if no jobs are assigned to them. Constraints (2) activate the presence of the batches if at least one job is assigned to them. Constraints (2)-(2) ensure that if a batch is used, its size is within the specified limits of the batch family. Constraints (2) guarantee that batches are scheduled once all the jobs in them are available. Constraints (2) schedules each batch on exactly one machine, if present. Finally, constraints (2) use the global constraint noOverlap, which ensures that the present interval variables in the provided sequence do not overlap. Hence, constraints (2) ensure that only one batch is processed at a time on each machine.
Figure 2 presents the optimal solution of the AS model that has a TWCT of $1,700$. The increase in the TWCT from the optimal (but infeasible) solution of the NS model is the result of a more restrictive model that enforces correctly assigned jobs to no overlapping batches. In the optimal solution of the problem, Jobs 1 and 3 are assigned to Batch 1, while Jobs 2 and 4 are assigned to Batch 2, and the batch interval variables do not overlap on the machine.
Figure 2: Optimal solution of the Assign-and-Schedule model
Although the beauty of the AS model is its clear distinction between assignment decisions with binary variables and scheduling decisions with interval variables and global constraints, this separation reduces the ability of the CP solver to prune the domains of the scheduling variables. The noOverlap constraints, in charge of handling the scheduling logic, provide the CP engine with a global view of the disjunctive constraints and effectively reduce the search space. However, constraints (2)-(2), in charge of handling the assignment while respecting the batch limits, are local in nature, reduce the overall propagation, and provide weak information to the global scheduling constraints. To remedy this, the models presented next make heavy use of global constraints for both the assignment and scheduling components.
VI-B The Synchronized Model
Model 3 The Synchronized Model Inspired by [10]
Variables:
$x_{jm}^{b}$$\in\{[a,a+p_{j})\}\cup\{\perp\},$ $\forall\leavevmode\nobreak\ j\in\mathcal{J},m\in\mathcal{M},b\in\mathcal{B}_{f_{j}};$ $(22)$
$y_{bm}$$\in\{[a,a+p^{f^{b}})\}\cup\{\perp\},$ $\forall\leavevmode\nobreak\ b\in\mathcal{B},m\in\mathcal{M};$ $(23)$
$\varphi_{m}$ $\in\text{Perm}\left(\set{y_{bm}}_{b\in\mathcal{B}}\right),$ $\forall\leavevmode\nobreak\ m\in\mathcal{M}.$ $(24)$
Functions:
${\mathscr{l}}_{m}^{b}$ $=\sum_{j\in\mathcal{J}_{f^{b}}}\text{pulse}\left(x^{b}_{jm},s_{j}\right),$ $\forall\leavevmode\nobreak\ m\in\mathcal{M},b\in\mathcal{B}.$ $(25)$
Formulation:
minimize $\text{minimize\leavevmode\nobreak\ }\sum_{j\in\mathcal{J}}\omega_{j}\cdot\sum_{b\in\mathcal{B}_{f_{j}}}\sum_{m\in\mathcal{M}}\texttt{endOf}\left(x^{b}_{jm}\right),$ $(26)$
subject to,
$\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\leavevmode\nobreak\ j\in\mathcal{J}:$ (27)
$\sum_{b\in\mathcal{B}_{f_{j}}}\sum_{m\in\mathcal{M}}\texttt{presenceOf}\left(x_{jm}^{b}\right)=1;$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\leavevmode\nobreak\ b\in\mathcal{B}:$ (28)
$\sum_{m\in\mathcal{M}}\texttt{presenceOf}\left(y_{bm}\right)\leq 1;$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\leavevmode\nobreak\ j\in\mathcal{J},b\in\mathcal{B}_{f_{j}},m\in\mathcal{M}$
synchronize $\left(x^{b}_{jm},\{y_{bm}\}\right);$ $(29)$
presenceOf $\left(x^{b}_{jm}\right)\Rightarrow\texttt{presenceOf}\left(y_{bm}\right);$ $(30)$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\leavevmode\nobreak\ m\in\mathcal{M}:$ (31)
noOverlap $\left(\varphi_{m}\right);$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\leavevmode\nobreak\ j\in\mathcal{J},m\in\mathcal{M},b\in\mathcal{B}_{f_{j}}:$ (32)
alwaysIn $\bigl{(}{\mathscr{l}}_{m}^{b},x_{jm}^{b},l_{f^{b}},u_{f^{b}}\bigr{)}.$
Model 3 presents the Synchronized (S) model. It is based on the CP model by Cakici et al. [10] for AM processes, which deals with a single compatible family; this paper extends their formulation to handle multiple incompatible families and include the necessary constraints to ensure a consistent synchronization of the jobs. Equation (3) defines an optional interval variable $x_{jm}^{b}$, of size $p_{j}$, that represents the option of job $j$ being batched in batch $b\in\mathcal{B}_{f_{j}}$ and processed on machine $m\in\mathcal{M}$. Equation (3) defines an optional interval variable $y_{bm}$ of size $p^{f^{b}}$ that represents the option of batch $b\in\mathcal{B}$ being processed on machine $m$. Equation (3) defines an interval sequence variable $\varphi_{m}$ that represents the sequence of batches scheduled on machine $m$. Equation (3) defines a cumulative function ${\mathscr{l}}_{m}^{b}$ that represents the load on machine $m$ due to the jobs in batch $b$.
Objective function (3) minimizes the TWCT. Constraints (3) assign each job to exactly one batch on exactly one machine. Constraints (3) allow each batch to be processed in at most one machine. Constraints (3) use the synchronize global constraint, which takes two inputs: an interval variable $v=x_{jm}^{b}$ and a set of interval variables $B=\set{y_{bm}}$. If $v$ is present, this constraint attempts to synchronize it with any intervals in $B$ that are also present [38], i.e., $\texttt{presenceOf}(v)=1\implies\texttt{startOf}(v)=\texttt{startOf}(b)\wedge\texttt{endOf}(v)=\texttt{endOf}(b),\forall\leavevmode\nobreak\ b\in B$ s.t. $\texttt{presenceOf}(b)=1$. In practice, when the interval variable representing the job being assigned to a batch on a machine (i.e., $x_{jm}^{b}$) is part of the solution, the synchronize constraint attempts to align its start and end times with those of the batch on the machine (i.e., $y_{bm}$), if it is also present. However, the synchronize constraint only ensures synchronization if both interval variables are present; it does not inherently enforce the presence of $y_{bm}$ when $x_{jm}^{b}$ is scheduled. This subtle issue, where the presence of $x_{jm}^{b}$ should logically imply the presence of $y_{bm}$, is not addressed in the original model by Cakici et al. [10], leading to synchronization inconsistencies. To resolve this, constraints (3) introduce an explicit implication constraint, which ensures that, whenever $x_{jm}^{b}$ is included in the solution, the corresponding batch interval $y_{bm}$ is also present, thereby guaranteeing consistent synchronization as intended. Constraints (3) use the noOverlap global constraint to ensure that the present interval variables in the machine sequence do not overlap, guaranteeing at most one batch being processed at a time on each machine. Finally, constraints (3) guarantee that the load of the machines when processing each job in its batch is within the size limits of the batch’s family.
Figure 3 shows the optimal solution of the S model, which has the same TWCT of $1,700$ as the AS model. Note how the S model enforces all the jobs in the same batch to start and end together in a synchronized fashion. In the context of the diffusion area in a wafer fab, this solution allows the thermal stability required by the finished products.
Figure 3: Optimal solution of the Synchronized model
The S model inspired by Cakici et al. [10] replaces the binary variables of the AS model with interval variables and uses more global constraints, like the synchronize and alwaysIn constraints. These changes make the S model more efficient than the AS model, as evidenced in the computational experiments. Nonetheless, some constraints of the S model still handle the assignment of jobs to batches through the binary presence variables of those intervals. This lack of global constraints to deal with the assignment in the model by Cakici et al. [10] leads to a weak propagation between the assignment and scheduling parts of the model. The novel CP model presented next exclusively uses global constraints to enhance computational performance.
VI-C The Redundant Synchronized Model
Model 4 presents the Redundant Synchronized (RS) model. It exclusively uses global constraints that exploit the problem structure to better guide the search process. Furthermore, it includes redundant variables and global constraints that increase the computational performance.
Equation (4) defines an interval variable $x_{j}$ of size $p_{j}$ that represents job $j\in\mathcal{J}$. Equation (4) defines an optional interval variable $x_{jm}$ that represents the option of processing job $j$ on machine $m\in\mathcal{M}$. Equation (4) defines an optional interval variable $x_{jm}^{b}$, also of size $p_{j}$, that represents the option of job $j$ being batched in batch $b\in\mathcal{B}_{f_{j}}$ and processed on machine $m$. Equation (4) defines an optional interval variable $y_{b}$ of size $p^{f^{b}}$ that represents batch $b\in\mathcal{B}$. Equation (4) defines an optional interval variable $y_{bm}$ of size $p^{f^{b}}$ that represents the option of batch $b$ being processed on machine $m$. Equation (4) defines an interval sequence variable $\varphi_{m}$ that represents the sequence of batches scheduled on machine $m$. Equation (4) defines the set $V_{jm}^{b}$ of interval variables that have to be synchronized with interval $x_{jm}^{b}$, if it is present. Equation (4) defines the unique state ${\mathfrak{f}}_{m}$ of the machine $m$ at every point in time. Equation (4) defines a cumulative function ${\mathscr{l}}_{mf}$ that represents the load on machine $m$ with jobs of family $f$.
Model 4 Redundant Synchronized model
Variables:
$x_{j}$$\in\{[a,a+p_{j})\},$ $\forall\leavevmode\nobreak\ j\in\mathcal{J};$ $(33)$
$x_{jm}$$\in\{[a,a+p_{j})\}\cup\{\perp\},$ $\forall\leavevmode\nobreak\ j\in\mathcal{J},m\in\mathcal{M};$ $(34)$
$x_{jm}^{b}$$\in\{[a,a+p_{j})\}\cup\{\perp\},$ $\forall\leavevmode\nobreak\ j\in\mathcal{J},m\in\mathcal{M},b\in\mathcal{B}_{f_{j}};$ $(35)$
$y_{b}$$\in\{[a,a+p^{f^{b}})\}\cup\{\perp\},$ $\forall\leavevmode\nobreak\ b\in\mathcal{B};$ $(36)$
$y_{bm}$$\in\{[a,a+p^{f^{b}})\}\cup\{\perp\},$ $\forall\leavevmode\nobreak\ b\in\mathcal{B},m\in\mathcal{M};$ $(37)$
$\varphi_{m}$ $\in\text{Perm}\left(\set{y_{bm}}_{b\in\mathcal{B}}\right),$ $\forall\leavevmode\nobreak\ m\in\mathcal{M}.$ $(38)$
Sets of variables to synchronize with $x_{jm}^{b}$:
$V_{jm}^{b}$ =$=\{y_{b},y_{bm},x_{jm}\},$ $\forall\leavevmode\nobreak\ j\in\mathcal{J},m\in\mathcal{M},b\in\mathcal{B}_{f_{j}};$ $(39)$
Functions:
[…]
${\mathscr{l}}_{mf}$ $=\sum_{j\in\mathcal{J}_{f}}\text{pulse}\left(x_{jm},s_{j}\right),$ $\forall\leavevmode\nobreak\ m\in\mathcal{M},f\in\mathcal{F}.$ $(41)$
Formulation:
minimize $\text{minimize \leavevmode\nobreak\ }\sum_{j\in\mathcal{J}}\omega_{j}\cdot\texttt{endOf}\left(x_{j}\right),$ $(42)$
subject to,
$\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\leavevmode\nobreak\ j\in\mathcal{J}:$ (43)
alternative $\left(x_{j},\set{x_{jm}}_{m\in\mathcal{M}}\right);$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\leavevmode\nobreak\ j\in\mathcal{J},m\in\mathcal{M}:$ (44)
alternative $\Bigl{(}x_{jm},\set{x^{b}_{jm}}_{b\in\mathcal{B}_{f_{j}}}\Bigr{)};$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\leavevmode\nobreak\ b\in\mathcal{B}:$ (45)
alternative $\left(y_{b},\set{y_{bm}}_{m\in\mathcal{M}}\right);$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\leavevmode\nobreak\ j\in\mathcal{J},b\in\mathcal{B}_{f_{j}},m\in\mathcal{M},v\in V_{jm}^{b}:$
presenceOf $\left(x^{b}_{jm}\right)\Rightarrow\texttt{presenceOf}\left(v\right);$ $(46)$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\leavevmode\nobreak\ j\in\mathcal{J},b\in\mathcal{B}_{f_{j}},m\in\mathcal{M}:$
synchronize $\left(x^{b}_{jm},V^{b}_{jm}\right);$ $(47)$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\leavevmode\nobreak\ m\in\mathcal{M}:$ (48)
noOverlap $\left(\varphi_{m}\right);$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\leavevmode\nobreak\ j\in\mathcal{J},m\in\mathcal{M}:$ (49)
alwaysEqual $\bigl{(}{\mathfrak{f}}_{m},x_{jm},f_{j}\bigr{)};$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\leavevmode\nobreak\ m\in\mathcal{M},f\in\mathcal{F},j\in\mathcal{J}_{f_{j}}:$ (50)
alwaysIn $\bigl{(}{\mathscr{l}}_{mf},x_{jm},l_{f},u_{f}\bigr{)}.$
Objective function (4) minimizes the TWCT. Constraints (4) ensure that each job is processed on exactly one machine. This can be seen as an assignment of jobs to machines, but now using the global alternative constraint, which exploits the structure of the problem to help the search process. Similarly, constraints (4) ensure that if a job is processed on a given machine, it is processed on such machine in exactly one batch. This can be seen as another assignment of jobs to batches on specific machines, but now using CP scheduling logic. Constraints (4) guarantee that each batch is processed on exactly one machine. This can be seen as another assignment of batches to machines with global constraints. The previous three constraint groups are completely independent and allow the machine assigned to each job, the batch where it is assigned, and the machine assigned to the batch, to be completely unrelated. Constraints (4) remedy this and link them all together, ensuring that if the interval variable $x^{b}_{jm}$ for the job $j\in\mathcal{J}$ in the batch $b\in\mathcal{B}_{f_{j}}$ on the machine $m\in\mathcal{M}$ is present in the final solution, then it must be the case that all the intervals in $V^{b}_{jm}$ must be present in the solution as well (i.e., the batch interval $y_{b}$, the interval of the batch on the machine $y_{bm}$, and the interval of the job on the machine $x_{jm}$). Furthermore, constraints (4) synchronize the start and end of the interval $x^{b}_{jm}$ with the intervals in $V^{b}_{jm}$, if they are present in the final solution. Constraints (4) ensure that only one batch is processed at a time on each machine. Constraints (4) ensure that jobs of only one family are processed at any time on the machines. Finally, constraints (4) guarantee that the load on the machines is within the required batch size limits of the family being processed.
VI-D Breaking symmetries
Symmetries in the search space could delay the search process to prove optimality. Hence, breaking these symmetries could potentially benefit the search process. Since the number of the batch is irrelevant for the solution, the following symmetry-breaking constraints, borrowed from Nascimento et al. [8], can be included in the AS and RS models.
$\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\leavevmode\nobreak\ f\in\mathcal{F},b\in\mathcal{B}_{f}\setminus\set{\min\mathcal{B}_{f}}:$
presenceOf $\left(y_{b}\right)\Rightarrow\texttt{presenceOf}\left(y_{b-1}\right);$ $(51)$
[…]
Constraints (VI-D) ensure that the batches of a family are used incrementally, i.e., if a batch $b$ is used, is because all the previous batches have already been used. Constraints (VI-D) go further by ensuring that the start times of the batches also increase with the batches. They use the global constraint startBeforeStart, which ensures that the start time of the first interval variable provided is before the start time of the second interval provided.
VII Computational experiments and results
Table II: The Structure of the instances.
Category Levels Definitions
Jobs 4 $|\mathcal{J}|\in\set{15,25,50,100}$
Families 2 ||$|\mathcal{F}|\in\set{3,5}$
Machines 2 ||$|\mathcal{M}|\in\set{2,3}$
Capacity 1 $u_{f}=50,l_{f}=1,\forall\leavevmode\nobreak\ f\in\mathcal{F}$
Processing times 2 $p^{f}\sim U\set{[\varsigma]},\forall\leavevmode\nobreak\ \varsigma\in\set{5,10}$
Job sizes 2 $s_{j}\sim U\set{[\varsigma]},\forall\leavevmode\nobreak\ \varsigma\in\set{25,50}$
Job weights 2 $\omega_{j}\sim U\set{[\varsigma]},\forall\leavevmode\nobreak\ \varsigma\in\set{5,10}$
Release time multipliers 2 $\lambda\in\set{0.5,1}$
# of instance classes 256
# of instances per class 10
Total # of instances 2,560
The computational experiments consider 2,560 instances following the same process used by Cakici et al. [6] and Ham et al. [18]. Table II presents the structure of the instances. The number of jobs in each instance can be one of four possible values: 15, 25, 50, or 100. The number of families in each instance can be one of two possible values: 3 or 5. The number of machines can be one of two possible values: 2 or 3. The size of the batches of each family must be between 1 and 50 in every instance. The processing times, job sizes, and job weights in each instance follow the distribution $U\set{[\varsigma]}$, which is the discrete uniform distribution over the values of the discrete set $[\varsigma]=\set{1,2,3,\ldots,\varsigma}$, where $\varsigma$ can be either 5 or 10 for the processing times and job weights, or either 25 or 50 for the job sizes. To generate the release times of the jobs, a lower bound of the overall makespan is computed as $C_{max,lb}=\left(\sum_{f\in\mathcal{F}}p^{f}\cdot\lceil N_{f}/u_{f}\rceil\right)/{|\mathcal{M}|}$. Then, this $C_{max,lb}$ is multiplied by $\lambda$ to define the possible set of release times. Hence $r_{j}\sim U\set{[\lambda\cdot C_{max,lb}]}$, where the $\lambda$ multiplier can take one of two possible values: 0.5 or 1. Overall, a total of $256=4\times 2^{6}$ instance classes result from the combination of the values and distributions described above. For each instance class (i.e., combination), 10 random instances were generated with batch size limits from 1 to 50, resulting in a grand total of $2,560$ instances.
The results only compare models that produce feasible solutions. Hence, each one of the generated instances was solved using the following six models:
  • NS+FR: Non-synchronized model by Ham et al. [18] + Feasibility Restoration step described in §V-B.
  • AS: Assign-and-Schedule model inspired by Nascimento et al. [8, 9].
  • AS+SB: AS model + Symmetry-breaking constraints.
  • S: Synchronized model inspired by Cakici et al. [10].
  • RS: The novel Redundant Synchronized model.
  • RS+SB: RS model + Symmetry-breaking constraints.
All the models and FR steps were implemented in Java 8. All the CP models were solved with IBM ILOG CP Optimizer [37] from CPLEX 22.1.1, using its Java interface [38]. All the experiments were run on the PACE Phoenix cluster [39], using machines that run Red Hat Enterprise Linux Server release 7.9 (Maipo) with dual Intel®  Xeon®  Gold 6226 CPU @ 2.70GHz processors, with 24 cores, and 48 GB RAM. Each run uses 8 cores and 16 GB RAM, and is given a time limit of 10 minutes (600 seconds) because a diffusion scheduling system in the semiconductor industry is expected to generate a Gantt-chart schedule every few minutes [18].
Figures 4-6 summarize the results. They group the 2,560 instances in 16 instance classes of 160 instances each, according to their number of jobs, families, and machines. Figure 4 presents the percentage of instances solved to optimality by each model in the top part of the chart, and the percentage of instances where each model could not find a solution in the bottom part. The key takeaway is that the RS model dominates all the other models by solving to optimality the most amount of instances with 15 and 25 jobs, and finding solutions to all the instances with 50 and 100 jobs. No model is able to prove optimality in the larger instances with 50 and 100 jobs. Furthermore, the AS and AS+SB models cannot even find solutions to some of these larger instances.
Figure 4: Percentage of instances solved to optimality (top) and with no solution found (bottom).
Figure 5 presents the average solve times obtained by each model. The stripes around the lines indicate the 95% confidence intervals. On the large instances with 50 and 100 jobs, most of the models reach the time limit given, which explains why none proves optimality in Figure 4. The NS+FR step is the fastest of all models, since the NS model produces infeasible solutions fast, and the FR step recovers their feasibility quickly. Nonetheless, as seen in Figure 4, this approach does not solve to optimality many instances compared to our RS model, which is the second fastest model.
Figure 5: Average solve time (s) by model with 95% confidence intervals.
To gain more insight on the quality of each model, the minimum TWCT is obtained across all models and is denoted by $\text{TWCT}_{i}^{*}$. Then, for each model, the gap on the instance as $|\text{TWCT}_{i,model}-\text{TWCT}_{i}^{*}|/|\text{TWCT}_{i,model}|$ is computed for each model. Figure 6 shows the average gap of each model on the instances where all the models found a solution. The stripes around the lines are the 95% confidence intervals. Although the NS+FR model is fast, its solutions reach up to 10% gap in the larger instances. It is particularly bad in instances with more machines since the FR step recovers the feasibility independently on each machine, being myopic to possible improvements by swapping jobs to other machines. The AS and AS+SB models perform better than the NS+FR model in the smaller instances but slightly worse in the larger ones. This is due to the lack of global constraints that help the search process. The S model performs slightly better than the AS model in the larger instances due to the replacement of the binary variables with interval variables and the addition of more global constraints. From the graph, it becomes clear that the RS model dominates all the other models, providing the best solutions overall. This is due to the redundancy in the structure of the problem and the exclusive use of global constraints. Interestingly, the SB constraints deteriorate the quality of the solutions obtained by the AS and RS models.
Figure 6: Average gap by model with 95% confidence intervals.
VIII Concluding remarks
This paper addressed the incompatible case of parallel batch scheduling, where compatible jobs belong to the same family, and jobs of different families cannot be processed together in the same batch. In the context of the diffusion area of the SM process, batches cannot be interrupted mid-process and jobs must be synchronized within their batch. The NS model by Ham et al. [18] for this problem does not synchronize the jobs within their batch, allowing interruptions mid-process. To remedy this limitation, the paper presented three new CP models, including two adaptions of existing models and a novel RS model. The AS model of Nascimento et al. [8] was extended to handle multiple incompatible job families and satisfy the batch size requirements. The S model of Cakici et al. [10] was also extended to the incompatible case while ensuring a consistent synchronization of jobs within their batches. Finally, the paper proposed a new RS model that exclusively uses global constraints and adds redundancy to the problem structure. The computational experiments demonstrate that the proposed RS model dominates all the other models: it finds solutions to all the instances, proves optimality in the largest amount of instances, produces solutions that are significantly better than other models, especially in the largest instances and achieves the best gaps in the least amount of time. Current research includes using CP for serial batching, and future research includes extending the efficient arc-flow formulations for compatible p-batch [16, 29, 28, 30] to the incompatible case and embedding the proposed formulations in larger scheduling contexts [40, 41, 42].
Acknowledgements
The authors would like to express their gratitude to Amira Hijazi for her support and helpful discussions.